there were 324 people at the school miusic featival. student tickets cost$2.50 and asb rickets cost $1.25. the total receipts were $645. find the number of student tickets and asb tickets sold.
Answer by mananth(16949)

x Students tickets
y asb tickets
Total attended
1 x + 1 y = 324 .............1

2.50 x + 1.25 y = 645 .............2
Eliminate y
multiply (1)by -1.25
Multiply (2) by 1
-1.25 x -1.25 y = -405
2.50 x + 1.25 y = 645
Add the two equations
1.25 x = 240
/ 1.25
x = 192
plug value of x in (1)
x+y= 324
192+y=324
y = 324-192
y = 132
192 Students tickets
132 asb tickets
m.ananth@hotmail.ca